OpenStudy (anonymous):

sq root x - 2 + 5 = 6

terenzreignz (terenzreignz):

\[\Large \sqrt{x-2}+5 = 6\] this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I want to learn the steps

terenzreignz (terenzreignz):

Yup :) First, put everything OUTSIDE THE SQUARE ROOT SIGN to the right-side of the equation.... For instance, this +5 here... \[\Large \sqrt{x-1}\color{red}{+5}= 6\] How to do that?

terenzreignz (terenzreignz):

Crud, sorry, typo.... \[\Large \sqrt{x-2}\color{red}{+5}= 6\]

terenzreignz (terenzreignz):

So you can do that by subtracting 5 from both sides... \[\Large \sqrt{x-2} \ {+ \ 5}\color{red}{-5}= 6\color{red}{-5}\]

terenzreignz (terenzreignz):

Leaving you with... \[\Large \sqrt{x-2}=\color{red}1\] can you do it from here?

OpenStudy (anonymous):

No

terenzreignz (terenzreignz):

Well, the most natural thing to do would be to square both sides, as squaring cancels out the radical... let me explain... \[\Large (\sqrt{a})^2 = a\] Catch me so far?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Actually, no, that makes no sense

terenzreignz (terenzreignz):

If you square a radical, you just remove the radical.

terenzreignz (terenzreignz):

Or, if you prefer it worded this way, if you raise a radical to an exponent 2, the exponent and the radical would "disappear" or cancel each other out.

OpenStudy (anonymous):

okay:) ready for the next step please

terenzreignz (terenzreignz):

Okay, so squaring both sides.... \[\Large (\sqrt{x-2})^2 = 1^2\] Now, \(\Large 1^2 = 1\) so it simplifies into... \[\Large (\sqrt{x-2})^2 = 1\] What does the left side simplify into?

OpenStudy (anonymous):

x^2?

terenzreignz (terenzreignz):

No... I just told you, squaring would just remove the radical \(\Large \sqrt{\color{white}{skjassj}}\) sign

OpenStudy (anonymous):

So nothing?

terenzreignz (terenzreignz):

Okay... back to square one... If we square a radical, both the square (the 2 exponent) and the radical "disappear" Case in point \[\Large (\sqrt3)^2 = 3\]\[\Large (\sqrt{a+b+c})^2 = a+b+c\] and so on. So, to what does the left side of this equation simplify into... \[\Large (\sqrt{x-2})^2 = 1\] ?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

this makes no sense okay? it really doesn't

OpenStudy (anonymous):

im sorry

terenzreignz (terenzreignz):

Perhaps we need a more thorough explanation? :)

terenzreignz (terenzreignz):

okay, we have a radical expression... with something, anything, inside the radical... \[\Large \left.\sqrt{<something>}\right.\] If we square it... or raise it to an exponent 2, \[\Large \left(\sqrt{<something>}\right)^2\] Then the \(\large 2\) and the radical sign \(\Large \sqrt{\color{white}{text}}\) would cancel each other out, leaving you with whatever was inside the radical... \[\Large \left(\sqrt{<something>}\right)^2=<something>\]